OpenStudy (anonymous):

Sue used a probability simulator to roll a 6-sided number cube and flip a coin 100 times. The results are shown in the tables below: Number on the Cube Number of Times Rolled 1 18 2 25 3 12 4 32 5 3 6 10 Heads Tails 56 44 81 over 100 56 over 100 1400 over 10000 1100 over 10000

OpenStudy (anonymous):

heads is 56 tails is 44

OpenStudy (astrophysics):

Ok what are we looking for

OpenStudy (anonymous):

Using Sue's simulation, what is the probability of rolling a 2 on the number cube and the coin landing heads up?

OpenStudy (anonymous):

i put 81/100 but i feel like that is wrong haha

OpenStudy (astrophysics):

Ok, so the total number of times rolled is 100, correct? (18+25+12+32+3+10) = 100

OpenStudy (anonymous):

mhm

OpenStudy (astrophysics):

So we have \[\frac{ 25 }{ 100 }\] for the number of times 2 was rolled in a 100 rolls of dice

OpenStudy (astrophysics):

Now we find the probability of dice, 56+44 = 100 (total amount flipped) We get \[\frac{ 56 }{ 100 }\] for the possibility of getting heads right

OpenStudy (astrophysics):

Now all we have to do is \[\frac{ 25 }{ 100 } \times \frac{ 56 }{ 100 }\] that will give us our total probability

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

could you help me with one more

OpenStudy (astrophysics):

No problem :)

OpenStudy (anonymous):

?

OpenStudy (astrophysics):

I can try post new question